OpenStudy (anonymous):

solve for x (sqrt(1-x^2)/2) - ((x^2-x)/(2sqrt(1-x^2))=0

OpenStudy (anonymous):

multiply the numerator and denominator of the left handside by sqrt(1=x^2) =\[\frac{ \sqrt{1-x ^{2}} }{ 2 }\times \frac{ \sqrt{1-x ^{2}} }{ \sqrt{1-x ^{2}} }-\frac{ (x ^{2}-x )}{ 2\sqrt{1-x ^{2}} }\]=0 we get =\[\frac{ (1-x ^{2})-(x ^{2}-x) }{ 2\sqrt{1-x ^{2}} }\]=0 we get \[-2x ^{2}+x+1=0\] is x=-1/2,1 ? yes or no?